# A Collaborative, Human-Centred Taxonomy of AI, Algorithmic, and Automation Harms

### v1

@@ Abstract.
This paper introduces a collaborative, human-centered taxonomy of AI, algorithmic and automation harms. We argue that existing taxonomies, while valuable, can be narrow, unclear, typically cater to practitioners and government, and often overlook the needs of the wider public. Drawing on existing taxonomies and a large repository of documented incidents, we propose a taxonomy that is clear and understandable to a broad set of audiences, as well as being flexible, extensible, and interoperable. Through iterative refinement with topic experts and crowdsourced annotation testing, we propose a taxonomy that can serve as a powerful tool for civil society organisations, educators, policymakers, product teams and the general public. By fostering a greater understanding of the real-world harms of AI and related technologies, we aim to increase understanding, empower NGOs and individuals to identify and report violations, inform policy discussions, and encourage responsible technology development and deployment.
@@ 1. Introduction
This paper introduces a collaborative and human-centered harms taxonomy aimed to serve as a powerful tool for civil society organisations, educators, policymakers, and product teams. We draw on other taxonomies, topic experts, and the 1,500+1limit-from5001,500+1 , 500 + incidents and issues captured in the AIAAIC Repository [^7] https://www.aiaaic.org/aiaaic-repository to set out a taxonomy of harms experienced across a wide variety of domains, countries, and cultures due to the inappropriate use and misuse of AI and adjacent technologies.
@@ 2. Related Work
Shelby et al. (2023) Sociotechnical harms AI community Peer-reviewed Survey on 172 papers
@@ 3.1. Working Group Setup And Core Aims
This taxonomy is the result of the collective effort of an independent working group. Since June 2023, 25+ volunteers from various backgrounds, nationalities, genders and levels of seniority, replied to an open call for participation, all of whom taking part in a personal capacity to develop a general purpose, outside-in, open taxonomy of near-term risks and harms of AI, algorithms, and automation [^13] List of organisations involved are listed in Appendix.
@@
  • Human-centered, based on actual incidents and their impact on different existing stakeholders
@@ 3.2.2. Literature review:
This initial mapping, along with a review of current literature exploring the risks and harms of AI and related technologies (Council, 2023; Hendrycks et al., 2023; Bengio et al., 2023; Tobin, 2023), informed the elaboration of a first version of the taxonomy, deliberately focused on near-term risks and documented harms of AI technologies. Key human rights and civil liberties charters, including the UN Universal Declaration of Human Rights, the European Convention on Human Rights, the Charter of Fundamental Rights of the European Union as well as the European Declaration on Digital Rights and Principles for the Digital Decade were also taken into account to produce the initial definition of the harms.
@@ 3.2.4. Annotation testing:
To facilitate this process, a dedicated open-source tool was developed by the working group. This user-friendly interface enables quick annotation through drop-down menus with reminders of the definitions, streamlined comment submission, faster human review and consensus assessment. This tool also generates, for each annotated incident, a Sankey diagram, visually depicting the annotators’ answers for each incident and computes its Krippendorff’s alpha coefficient, a score measuring agreement–or lack thereof–among annotators (Gwet, 2014) [^16] It is expected that the average alpha coefficient go up and progressively tend towards 1 -perfect agreement- as the taxonomy is improved, while never reaching it (standard deviation due to crowd-sourced annotations, emergence of new use cases…), providing the working group with a clear KPI., allowing for a swift and objective identification of potential gaps, classification conflicts or contentious definitions. During the 9 first rounds of annotation, from February to April 2024, the working group generated over 1,000 individual annotations on 39 incidents pulled from the AIAAIC Repository.
@@ 4. Overview of the Harms Taxonomy
For the purposes of this taxonomy, a harm is defined as a “physical, pyschological or other form of damage to third-parties” that results from the use of an AI, algorithmic or automation system. The taxonomy organises harms at two different levels of granularity. The first level is the “harm type”, which describes the general category of harm. The second level divides harm types into sub-categories, referred to as “specific harms”. Definitions are provided for each harm category and specific harm.
@@
To facilitate a comparative understanding of our proposed taxonomy, we have selected other prominent taxonomies based on their scope, purpose, and intended audience, ensuring a relevant and meaningful comparison. The comparison, outlined in Table 2, features frameworks developed by recognised institutions such as the OECD, EPIC, CSET, the Alan Turing Institute, Microsoft Azure, Sony, and SHAS (Rismani et al., 2023). These taxonomies were chosen because they similarly address a broad range of AI-related harms and are designed for diverse uses including policy guidance, risk management, and public education, and are intended for a wide audience encompassing policymakers, industry stakeholders, and the general public. Each column in the table represents a different taxonomy and indicates whether specific types of harm are addressed within their frameworks, highlighting how the AI harms taxonomy aligns with or differs from established approaches in the field. Among the taxonomies we compared, those developed by the OECD, CSET, and SHAS are most similar to ours due to their comprehensive scopes and focus on both policy-oriented applications and broad societal impacts. These frameworks emphasise a wide coverage of harms, both tangible and intangible, aligning closely with our objective to inform risk management and policy guidance comprehensively. The SHAS taxonomy, in particular, shares our emphasis on sociotechnical aspects, making it highly relevant for interdisciplinary approaches to understanding and mitigating AI-related harms.
@@ 5. Benefits and Intended Uses
Such models should be based on both the likelihood and the expected harm of an incident. While the former can be measured on the basis of statistical metrics that measure the compliance of AI to a set of Safe and trustworthy requirements (see e.g. (Babaei et al., 2024) for a consistent set of such metrics), the latter require an assessment of all possible harms that non-compliance can lead to. Traditional taxonomies focus on the harms caused to a private or public organisation, or to a national entity; however, our taxonomy is able to capture a wide spectrum of harms, occurring to individuals, organisations, society and the environment, across the world, and is therefore better equipped to estimate risks, and prioritise consequent mitigation actions.
@@ 6. Conclusion and Future Work
Benefits: The development of this taxonomy is human centered and volunteer-based. The use of taxonomies to classify content, and in particular annotation has a dark side. There are examples of large technology companies exploiting gig workers for the completion of annotation tasks, including the classification of harmful content (Xiang, 2023). The taxonomy developed here, and to be used for classifying and annotating events, is human centered in the sense not just that it is intended to be understood and used by end users as well as experts, but also in the sense that it is based on volunteer input as opposed to an exploitative employment contract.
@@ A.8. Political and Economic
- Erosion of trust in public institutions and weakened checks and balances due to mis/disinformation, influence operations, overdependence on technology, etc.

### v2

@@ Abstract.
This paper introduces a collaborative, human-centred taxonomy of AI, algorithmic and automation harms. We argue that existing taxonomies, while valuable, can be narrow, unclear, typically cater to practitioners and government, and often overlook the needs of the wider public. Drawing on existing taxonomies and a large repository of documented incidents, we propose a taxonomy that is clear and understandable to a broad set of audiences, as well as being flexible, extensible, and interoperable. Through iterative refinement with topic experts and crowdsourced annotation testing, we propose a taxonomy that can serve as a powerful tool for civil society organisations, educators, policymakers, product teams and the general public. By fostering a greater understanding of the real-world harms of AI and related technologies, we aim to increase understanding, empower NGOs and individuals to identify and report violations, inform policy discussions, and encourage responsible technology development and deployment.
@@ 1. Introduction
This paper introduces a collaborative and human-centred harms taxonomy aimed to serve as a powerful tool for civil society organisations, educators, policymakers, and product teams. We draw on other taxonomies, topic experts, and the 1,500+1limit-from5001,500+1 , 500 + incidents and issues captured in the AIAAIC Repository [^7] https://www.aiaaic.org/aiaaic-repository to set out a taxonomy of harms experienced across a wide variety of domains, countries, and cultures due to the inappropriate use and misuse of AI and adjacent technologies.
@@ 2. Related Work
Shelby et al. (2023) Socio-technical harms AI community Peer-reviewed Survey on 172 papers
@@ 3.1. Working Group Setup And Core Aims
This taxonomy is the result of the collective effort of an independent working group. Since June 2023, 25+ volunteers from various backgrounds, nationalities, genders and levels of seniority, replied to an open call for participation, all of whom taking part in a personal capacity to develop a general purpose, outside-in, open taxonomy of near-term risks and harms of AI, algorithms, and automation [^13] List of individuals involved are listed in Acknowledgments.
@@
  • Human-centred, based on actual incidents and their impact on different existing stakeholders
@@ 3.2.2. Literature review:
This initial mapping, along with a review of current literature exploring the risks and harms of AI and related technologies (Norwegian Consumer Council, 2023; Hendrycks et al., 2023; Bengio et al., 2024; Tobin, 2023), informed the elaboration of a first version of the taxonomy, deliberately focused on near-term risks and documented harms of AI technologies. Key human rights and civil liberties charters, including the UN Universal Declaration of Human Rights, the European Convention on Human Rights, the Charter of Fundamental Rights of the European Union as well as the European Declaration on Digital Rights and Principles for the Digital Decade were also taken into account to produce the initial definition of the harms.
@@ 3.2.4. Annotation testing:
[…]
A screenshot of the annotation tool, showcasing the active ’Annotator’ section. The left pane provides navigation links to ’Instructions,’ ’Annotator,’ and ’Results,’ while the right pane displays a series of classification questions to be answered.
To facilitate this process, a dedicated open-source tool was developed by the working group. This user-friendly interface enables quick annotation through drop-down menus with reminders of the definitions, streamlined comment submission, faster human review and consensus assessment. This tool also generates, for each annotated incident, a Sankey diagram, visually depicting the annotators’ answers for each incident and computes its Krippendorff’s alpha coefficient, a score measuring agreement–or lack thereof–among annotators (Gwet, 2014) [^16] The average alpha coefficient is expected to increase progressively towards 1, indicating perfect agreement, as the taxonomy improves. However, it will likely never reach 1 due to standard deviation from crowdsourced annotations and the emergence of new use cases. This provides the working group with a clear KPI., allowing for a swift and objective identification of potential gaps, classification conflicts or contentious definitions. During the 9 first rounds of annotation, from February to April 2024, the working group generated over 1,000 individual annotations on 39 incidents pulled from the AIAAIC Repository.
@@ 4. Overview of the Harms Taxonomy
Figure 2. An overview of the AI, algorithmic and automation harms taxonomy. A printer-friendly version is available in fig. 3 in appendix A
[…]
@@
For the purposes of this taxonomy, a harm is defined as a “physical, psychological or other form of damage to third-parties” that results from the use of an AI, algorithmic or automation system. The taxonomy organises harms at two different levels of granularity. The first level is the “harm type”, which describes the general category of harm. The second level divides harm types into sub-categories, referred to as “specific harms”. Definitions are provided for each harm category and specific harm.
@@
To facilitate a comparative understanding of our proposed taxonomy, we have selected other prominent taxonomies based on their scope, purpose, and intended audience, ensuring a relevant and meaningful comparison. The comparison, outlined in Table 2, features frameworks developed by recognised institutions such as the OECD, EPIC, CSET, the Alan Turing Institute, Microsoft Azure, Sony, and SHAS (Rismani et al., 2023). These taxonomies were chosen because they similarly address a broad range of AI-related harms and are designed for diverse uses including policy guidance, risk management, and public education, and are intended for a wide audience encompassing policymakers, industry stakeholders, and the general public. Each column in the table represents a different taxonomy and indicates whether specific types of harm are addressed within their frameworks, highlighting how the AI harms taxonomy aligns with or differs from established approaches in the field. Among the taxonomies we compared, those developed by the OECD, CSET, and SHAS are most similar to ours due to their comprehensive scopes and focus on both policy-oriented applications and broad societal impacts. These frameworks emphasise a wide coverage of harms, both tangible and intangible, aligning closely with our objective to inform risk management and policy guidance comprehensively. The SHAS taxonomy, in particular, shares our emphasis on socio-technical aspects, making it highly relevant for interdisciplinary approaches to understanding and mitigating AI-related harms.
@@ 5. Benefits and Intended Uses
Such models should be based on both the likelihood and the expected harm of an incident. While the former can be measured on the basis of statistical metrics that measure the compliance of AI to a set of Safe and trustworthy requirements (see e.g. (Babaei et al., 2025) for a consistent set of such metrics), the latter require an assessment of all possible harms that non-compliance can lead to. Traditional taxonomies focus on the harms caused to a private or public organisation, or to a national entity; however, our taxonomy is able to capture a wide spectrum of harms, occurring to individuals, organisations, society and the environment, across the world, and is therefore better equipped to estimate risks, and prioritise consequent mitigation actions.
@@ 6. Conclusion and Future Work
Benefits: The development of this taxonomy is human centred and volunteer-based. The use of taxonomies to classify content, and in particular annotation has a dark side. There are examples of large technology companies exploiting gig workers for the completion of annotation tasks, including the classification of harmful content (Perrigo, 2023). The taxonomy developed here, and to be used for classifying and annotating events, is human centred in the sense not just that it is intended to be understood and used by end users as well as experts, but also in the sense that it is based on volunteer input as opposed to an exploitative employment contract.
@@ Appendix A Appendix A
For the printer-friendly version of the taxonomy diagram, please refer to fig. 3 at the end of the document.
@@ A.8. Political and Economic
- Erosion of trust in public institutions and weakened checks and balances due to mis/disinformation, influence operations, over-dependence on technology, etc.
@@ A.9. Environmental
[…]
A tree diagram illustrating the hierarchical structure of the AI, algorithmic, and automation harms taxonomy. It displays the main categories and their respective subcategories.